OpenStudy (anonymous):

Write in point-slope form an equation of the line through each pair of points. (-2, 3) and (2, 9)

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

y-y1=m(x-x1) y-3=m(x-2)

OpenStudy (anonymous):

y-9=m(x-2)

OpenStudy (anonymous):

You need to find the gradient use the formula: